# Computational complexity of the recoverable robust shortest path problem in acyclic digraphs

### v1

@@ 2 Computational complexity results
Let $G=(V,A)$, $|V|=n$, $|A|=m$, be an instance of Hamiltonian Path. We build the corresponding instance of Rec Rob SP as follows. The graph $G^{\prime}$ has $2n$ copies of the nodes in $V=\{v_{1},\dots,v_{n}\}$, labeled as $v_{i}^{j}$, $i\in[n]$, $j\in[2n]$, and two additional nodes denoted by $s$ and $t$. The arcs of $G^{\prime}$ are formed as follows. The vertical arcs connect $v_{i}^{j}$ with $v_{i}^{j+1}$ for $i\in[n]$, $j\in[2n-1]$, node $s$ with $v_{i}^{1}$, $i\in[n]$, and nodes $v_{i}^{2n}$ with $t$, $i\in[n]$. The vertical arcs form $n$ disjoint vertical $s$-$t$ paths $P_{1},\dots,P_{n}$ in $G^{\prime}$, all of which have exactly $12n+$ 1 arcs. The diagonal arcs connect $v_{i}^{2j}$ with $v_{k}^{2j+1}$ for $(v_{i},v_{k})\in A$, $j\in[n-1]$ (the graph $G^{\prime}$ is constructed in the same way as in [1, the proof of Theorem 2], where the min-max regret shortest path problem under the interval uncertainty is considered). A sample construction is shown in Figure 1. Set any positive budget $\Gamma^{c}>0$, say $\Gamma^{c}=1$, and the recovery parameter $k=2n$ for arc-inclusion (arc-exclusion neighborhood) and $k=4n$ for the arc-symmetric difference neighborhood. All arcs in $G^{\prime}$ have first-stage costs equal to 0. The second-stage costs of the vertical arcs are in the interval $[0,M]$, and the second-stage costs of the diagonal arcs are in the interval $[M,M]$, where $M$ is any constant greater than $\Gamma^{c}$. We will show that $G$ has a Hamiltonian path if and only if there is a solution to Rec Rob SP in $G^{\prime}$ with the total cost at most $\frac{\Gamma^{c}}{n}$.
Figure 1: A sample digraph $G$ and the corresponding digraph $G^{\prime}$.
$\Rightarrow$ Let $P$ be a Hamiltonian path in $G$ that visits the nodes of $G$ in order $v_{i_{1}},\dots,v_{i_{n}}$. Consider path $X$ in $G^{\prime}$ that contains the arcs
$(s,v_{i_{1}}^{1}),(v_{i_{1}}^{1},v_{i_{1}}^{2}),(v_{i_{1}}^{2},v_{i_{2}}^{3}),(v_{i_{2}}^{3},v_{i_{2}}^{4}),(v_{i_{2}}^{4},v_{i_{3}}^{5}),\dots(v_{i_{n}}^{2n},t)$
These arcs exist by the construction of $G^{\prime}$. An example is shown in Figure 1. The Hamiltonian path in $G$, which visits the nodes of $G$ in order $v_{1},v_{3},v_{2},v_{4}$, corresponds to the path composed of the arcs shown in red on the corresponding graph $G^{\prime}$. Observe that $X$ uses at least one arc from each vertical path $P_{1},\dots,P_{n}$. Therefore, each vertical path $P_{i}$ is in $\Phi(X,k)$ and can be used as the second-stage (recovery) path. In the worst second-stage cost scenario for $X$, the budget $\Gamma^{c}$ is shared equally among the arcs of all vertical paths (all vertical paths must have the same second-stage cost). Because the first-stage cost of $X$ is 0 and the second-stage cost of the best second-stage path is $\frac{\Gamma^{c}}{n}$, the total cost of the solution to Rec Rob SP is $\frac{\Gamma^{c}}{n}$.
$\Leftarrow$ Assume that there is a solution to Rec Rob SP with a total cost of at most $\frac{\Gamma^{c}}{n}$. Let $X$ be the first-stage path in this solution. We claim that $X$ must use at least one arc from each vertical path $P_{1},\dots,P_{n}$. Indeed, suppose that one of these paths, say $P_{l}$, does not intersect with $X$. Because $k=2n$, the path $P_{l}\notin\Phi(X,k)$. Since the second-stage costs of the diagonal arcs are equal to $M\geq\Gamma^{c}$, no diagonal arc can be used in the second-stage path. Therefore, in a worst scenario for $X$, the budget $\Gamma^{c}$ is not allocated to the arcs in path $P_{l}$. Consider scenario $\boldsymbol{c}$ in which $\Gamma^{c}$ is equally shared to all the arcs of the remaining $n-1$ vertical paths. In what follows, every second-stage path in $\Phi(X,k)$ has a cost of at least $\frac{\Gamma^{c}}{n-1}$, which contradicts the assumption that there is a solution to Rec Rob SP with the total cost of at most $\frac{\Gamma^{c}}{n}$. We can now retrieve a Hamiltonian path in $G$ from the diagonal arcs in $X$. Notice first that $X$ has exactly $12n+$ 1 arcs ($G^{\prime}$ is a layered digraph) and contains exactly $n-1$ diagonal arcs. Indeed, the diagonal arcs must connect $n$ disjoint vertical paths, so the number of diagonal arcs is at least $n-1$. On the other hand, the number of diagonal arcs cannot be greater than $n-1$, since the number of arcs in $X$ is equal to $12n+$ 1 (we need $n$ arcs from the vertical paths and 2 arcs connecting $s$ and $t$ in $X$). Now, it is easy to see that the diagonal arcs in $X$ correspond to the arcs on a Hamiltonian path in $G$. ∎
@@
[…]

### v2

@@ 2 Computational complexity results
The starting point is the gap-preserving reduction from Max 3-Sat to Rec Rob SP (see Lemma 2). By combining this reduction with the reduction from 3-Sat to Max 3-Sat (see Lemma 1) we get a reduction from 3-Sat to Rec Rob SP, which for some $\delta\in(0,1)$ (e.g., $\delta=1/(1+\gamma))$ that ensures:
  1. 1.
    if $\Psi$ is satisfiable then $OPT(\mathcal{I}^{(q)})=\frac{1}{m^{q}}$,
  2. 2.
    if $\Psi$ is not satisfiable then $OPT(\mathcal{I}^{(q)})>\frac{1}{\delta^{q}m^{q}}$.
It has been shown in [9] that $\delta=\frac{7}{8}$.
Let $q=(\log n)^{\beta}$ for some fixed $\beta>0$, where $n$ is the number of variables in an instance of 3-Sat. Now, the size of $\mathcal{I}^{(q)}$ is $N=n^{O(q)}=n^{O(\log^{\beta}n)}=2^{O(\log^{\beta+1}n)}$. Therefore, the running time of the reduction is quasi-polynomial $n^{\text{poly}(\log n)}$. The resulting gap of $\mathcal{I}^{(q)}$ is $\left(\frac{1}{\delta}\right)^{q}=\left(\frac{1}{\delta}\right)^{\log^{\beta}n}=2^{O(\log^{\beta/\beta+1}N)}$.
Assume, on the contrary, that there is a polynomial-time algorithm for Rec Rob SP problem with an approximation ratio of $2^{\log^{1-\epsilon}N}$ for any $\epsilon>1-\frac{\beta}{\beta+1}$. Applying this algorithm to the resulting instance $\mathcal{I}^{(q)}$, one could evaluate whether a formula $\Psi$ in 3-SAT is satisfiable in $O(n^{\mathrm{poly}\log n})$ time. But this would imply NP $\subseteq$ DTIME$(n^{\mathrm{poly}\log n})$, a contradiction. ∎
Corollary 1.
The Rec Rob SP problem in general acyclic digraphs with any neighborhood (1)-(3), under $\mathcal{U}(\Gamma^{c})$, is strongly NP-hard and not approximable within any constant $\alpha>0$, unless P=NP, if $k$ is part of the input.
Proof.
It is sufficient to set $q=\frac{\log\alpha}{\log 1/\delta}$ in the proof of Theorem 2. ∎
We now consider discrete budgeted uncertainty. First, we recall that the Rec Rob SP problem is weakly NP-hard even for a very restricted class of graphs.
Theorem 3 ([8]).
The Rec Rob SP problem is weakly NP-hard in arc series-parallel digraphs with any neighborhood (1)-(3), under $\mathcal{U}(\Gamma^{d})$, if $k$ is part of the input.
In the case of more general graphs, the problem becomes significantly more challenging.
@@
The Rec Rob SP problem is strongly NP-hard in layered digraphs with any neighborhood (1)-(3), under $\mathcal{U}(\Gamma^{d})$, and it is also not at all approximable, unless P=NP, if $k$ is part of the input.